# Implications of computer science theory for the simulation hypothesis

### v1

@@ Abstract
The simulation hypothesis has recently excited renewed interest, especially in the physics and philosophy communities. However, the hypothesis specifically concerns computers that simulate physical universes, which means that to properly investigate it we need to couple computer science theory with physics. Here I do this by exploiting the physical Church-Turing thesis. This allows me to introduce a preliminary investigation of some of the computer science theoretic aspects of the simulation hypothesis. In particular, building on Kleene’s second recursion theorem, I prove that it is mathematically possible for us to be in a simulation that is being run on a computer by us. In such a case, there would be two identical instances of us; the question of which of those is “really us” is meaningless. I also show how Rice’s theorem provides some interesting impossibility results concerning simulation and self-simulation; briefly describe the philosophical implications of fully homomorphic encryption for (self-)simulation; briefly investigate the graphical structure of universes simulating universes simulating universes, among other issues. I end by describing some of the possible avenues for future research that this preliminary investigation reveals.
@@ 1.1 Background
This modern iteration of the hypothesis has many versions that vary in their precise details. However, the central idea is that some portion of a physical universe, including some conscious reasoning agents that exist in that cosmological universe, might in fact be part of a simulation being run in a physical computer of some super-sophisticated race of aliens [12, 31, 20]. Under the simulation hypothesis, those agents would be variables evolving in a computer program running in a physical computer designed by such a putative super-sophisticated race of aliens. In particular, it might be that a portion of our cosmological universe is being simulated this way, so that we are variables evolving in some simulation computer. The vague implication is that if we are in fact a simulation, then we do not have any sort of “objective reality”. Moreover, much of philosophy of science, going back to Kant (at least), would be rendered moot if we are simulations. In particular, this would happen to many of the flavors of “*** realism”, which suppose there “exists” some “real” physical truth which is not reducible to a mathematical formalism, but instead is somehow “concrete”, non-mathematical in its very essence — and so not just a computer algorithm. [^1] Arguably, the simple fact that the simulation hypothesis is logically consistent, but if true would imply that there is no such “concrete”, non-mathematical reality, establishes that it is impossible to establish the logical necessity of any of these flavors of realism,
@@
These investigations of relative probabilities all beg (many) questions, about what such probabilities might actually mean, formally. (Just as there are such questions concerning the idea of ascribing probabilities to the universes in a level IV multiverse in general.) How does one ascribe a measure, obeying the Kolmogorov axioms based on some associated sigma algebra, to a collection of event each of which is a universe? Note that it is not even clear that this collection would be a set, rather than a proper class of some sort. Directly reflecting such problems, one could not experimentally assess any proposed value of such a probability, e.g., with a proper scoring rule.  [^2] One might try to use a Bayesian “degree of belief” interpretation of probability to circumvent this issue. However, without any decision that would be made based on the probability assigned, and associated loss function, one could not apply Bayesian decision theory. So it is hard to see how “degree of belief” is meaningful in this case. See [61] for some related arguments.
More prosaically, many papers considering such probabilities assume, implicitly or otherwise, that there is some way to assign a uniform probability to all universes. It is proven below in Section 2.4 that this is mathematically impossible in the kind of universes considered in this paper (which, one might argue, is the kind of universes considered in these other papers). One might imagine that rather than a uniform probability distribution, one could assign some sort of Cantor (“fair coin”) measure to the set of possible universes, as is done for example in algorithmic information theory. However, it is proven in Section 7 that it is impossible to assign a Cantor measure to a set of universes defined by a natural and interesting set of restrictions.
@@
Yet others have considered issues concerning the simulation hypothesis that might charitably be characterized as “philosophical” in nature [32, 21, 26], at least in that they do not involve mathematical reasoning of any sort.
@@
One earlier set of semi-formal results along these lines was presented in [64]. Some of those results define “simulation” in terms of the relationship between languages at different levels of the time, polynomial, or exponential hierarchies of CS theory, and some define “simulation” in terms of the relationship between languages at different levels the arithmetic or analytic hierarchies of logic, or similar constructions [22, 5, 30]. Other results in that paper define “simulation” in terms of the relationship between computational machines with different Turing degrees [57, 54]. Yet others consider the application of Gödel’s second incompleteness theorem, defining simulation in terms of languages (in the logic theory sense) with nested sets of axioms.
None of those results consider what it means for a computational machine to simulate a physical universe, per se. There is no concern for “coupling” the mathematics of CS theory to the laws of physics of our universe. One natural way to address this lacuna is to use a particular formalization of the physical Church-Turing thesis (PCT), and a closely related thesis which I call the reverse PCT (RPCT). That is the approach I adopt in this paper. [^3] The reader unfamiliar with TMs and associated concepts like Universal Turing machines (UTMs), instantaneous descriptions (ID) of the state of a TM and its tapes, prefix-free encodings, etc., should consult Appendix A for appropriate background. Specifically, in this paper I will say that the PCT applies to a particular (portion of a) universe if the dynamics of that (portion of a) universe can be implemented on a Turing machine (TM). The RPCT instead says that any desired TM can be implemented by the dynamics of the universe in question, by appropriate choice of the initial conditions of that universe.
@@ 2.1 Notation
Other important definitions and notations is in Appendix A , e.g., of partial functions, and of computable functions (here always assumed to be total). I also review the definition of “computational universality” in that appendix. In addition, there I review the definitions of a universal TM (UTM), and a prefix-free TM (the implementation of TMs I will often be assume in this paper, implicitly or otherwise). I also define the instantaneous description (ID) of a Turing machine (TM) there. In the main body of this paper I will assume that the non-blank alphabet of the TM, $\Sigma$, is just ${\mathbb{B}}$. So as described in Appendix A, we can take the state of the tape to be a finite string in ${\mathbb{B}}^{*}$, even though strictly speaking the definition of TMs assumes infinitely long tapes.
@@ 2.2 Framework for analyzing CS theory of the simulation hypothesis
where $W$ is initially finite and cannot be extended dynamically, whereas $N$ is countably infinite. I will parameterize $N$ as $\underline{N}\times R$. $\underline{N}={\mathbb{B}}^{*}$ is the state of the tape(s) in the case that the computer is represented as a TM, or it is the state of the memory in the case that the computer is represented as a RAM machine, or an appropriately modified laptop computer. The reason for writing the countably infinite space $\underline{N}$ as the set of finite bit strings is so that we can physically implement it as a system that initially has a finite state space, but whose state space can be expanded by an arbitrarily large finite amount an arbitrary number of times as the universe evolves. See Section 2.6. I will write the elements of $\underline{N}$ as $\underline{n}$.
$R$ instead is a finite set that represents the internal variables of the computer. So for example, in a conventionally represented TM, $R$ would be the state of the TM’s head, the position(s) of its pointer(s) on the tape(s), etc.  [^5] Strictly speaking, in a TM the pointer variables can have any value in ${\mathbb{N}}$, since they can point to an arbitrary position on the (countably infinite) tape(s). This technical concern can be addressed several ways. For example, we might want to have the states of the pointer variables at any iteration after the $k$’th iteration given by the state of some special subsection of the tape(s), $\underline{n}$, with the states of those variables at earlier iterations recorded in $R$. Another possibility would be to have $R$ infinitely extendable.
@@
To simplify the analysis, I do not assume that the models of universes that I investigate in this paper capture those universes in toto, e.g., if the state spaces of those entire universes are actually uncountably infinite. Similarly, I will not assume that a simulation running in a computer inside a universe models the dynamics of an entire universe. Rather throughout this paper the expression “universe” should be understand as shorthand for “possibly coarse-grained subset of a universe”. However I do assume that the universe obeys deterministic dynamics, be it over the original state space (assuming it is countable) or some coarse-grained state space. As an example, in Section 2.6 I present a detailed example of what such a “coarse-grained subset of a universe” could be for our particular physica universe.
@@ 2.3 What it means for one universe to simulate another
In this paper I will formalize simulation as follows:
@@
Note that the LHS of Eq. 5 is the second of the two components of the vector-valued function $g$, while the RHS is an encoding of both components of $g^{\prime}$ into a single variable. Note also the requirement in Definition 1 that the state of the simulating computer $N$ not change after it completes its simulation of the future state of $V^{\prime}$. In other words, I require that the computer halts when it completes its simulation. I also require that the ID of the computer $N$ be replaced in toto (i.e., uniquely, with all other variables fixed to some predefined values) by the output of its simulation program. So for example, if $N$ is a multi-tape prefix TM, this means that when $N$ halts with the result of its simulation on its output tape, all the other tapes — the intermediate work tapes and the input tape — have been re-initialized to be all blanks.
@@
Note that Definition 1 does not require that the simulating computer $N$ calculates the future state of the universe $V^{\prime}$ no matter what the initial state of the environment $W$ outside of $N$, i.e., no matter what the value of $w_{0}$. The reason for this is so that the computer $N$ can retrieve the specific information it needs to perform its simulation of the dynamics of the specific state $v^{\prime}_{0}$ from the super-aliens running $N$, who exist in the environment of $N$, $W$. Note though that this freedom also allows the beings running the simulation computer $N$ to intervene on the dynamics of that computer, e.g., by overwriting the simulation program being run in the computer at any time they want. They can even “pull the plug early” on that computer, before it finishes its computation.
@@
I will informally use the term cosmological universe to refer to an entire physical universe obeying one set of laws of physics throughout, with one set of shared initial conditions, etc. [^6] For simplicity, I sidestep the issue of multiverses with different physical constants but otherwise identical physical laws, e.g., arising from a shared inflation epoch. In general, $V$ and $V^{\prime}$ might be portions of different cosmological universes, obeying different laws of physics. They could also be parts of the same cosmological universe though (and therefore obeying the same laws of physics). One way this could occur is if there are physically distinct regions of same universe. (See Section 2.6.) This example of the situation will be important in the discussion of the simulation graph below in Section 7.
@@ 2.4 The Physical Church-Turing Thesis
It is also worth pointing out that some readers might argue that Definition 3 doesn’t exactly capture any of the various properties that have been referred to as the “physical Church Turing thesis” in the literature. In some senses it more like one of the related concepts inspired by modern physics, e.g., some forms of ontic structural realism [28, 39, 38, 4] or the level IV multiverse [58, 59]. But for current purposes, we can ignore these semantic distinctions.
@@ 2.5 The Reverse Physical Church-Turing Thesis
Broadly speaking, the RPCT says that the system $N$ operates like a UTM for all pairs $(k,y)$ such that $T^{k}(y)$ is defined, where $k$ and/or $y$ may be encoded in some degrees of freedom in $w_{0}$ rather than directly in $n_{0}$. (That freedom to have $k$ and / or $y$ in the environment of the computer allows that computer to do things like observe its environment to get the input for a computation that it then runs.)
@@
Eqs. 6 and 7 mean that the physical system implementing the UTM is initialized with the precise TM that it is to supposed to implement, but not the actual data that TM will be run on. In general, that data (the value of $y$) would be transferred into the computer $N$ at a subsequent iteration, after this initialization of the UTM, but before the UTM starts to run. As an example, this is what would occur if we were to initialize a laptop with a universe-simulating program, with the precise data that program is to run on fed into the laptop before starting the simulation program. (See Section 2.6.) In this paper, when assuming the RPCT I will not assume that the pristine RPCT holds in particular unless I explicitly say so.
@@
Nonetheless, even if the PCT holds in our specific universe, that does not mean that the RPCT has to hold. In particular, cosmological considerations could prevent it from holding [37]. So the results below specific to universes that obey the PCT and / or universes that obey the RPCT might not apply to our specific universe. However, approximations of the analysis might apply even if the PCT and/or RPCT don’t hold exactly in our universe, depending on how precisely they are violated.

### v2

@@ Abstract
The simulation hypothesis has recently excited renewed interest, especially in the physics and philosophy communities. However, the hypothesis specifically concerns computers that simulate physical universes, which means that to formally investigate it we need to couple computer science theory with physics. Here I couple those fields with the physical Church-Turing thesis. I then exploit that coupling to investigate of some of the computer science theory aspects of the simulation hypothesis. In particular, I use Kleene’s second recursion theorem to prove that it is mathematically possible for us to be a simulation that is being run on a computer — by us. In such a self-simulation, there would be two identical instances of us; the question of which of those is “really us” is meaningless. I also show how Rice’s theorem provides some interesting impossibility results concerning simulation and self-simulation; briefly describe the philosophical implications of fully homomorphic encryption for (self-)simulation; and briefly investigate the graphical structure of universes simulating universes simulating universes $\ldots$, among other issues. I end by describing some of the possible avenues for future research.
@@ 1.1 Background
This modern iteration of the hypothesis has many versions that vary in their precise details. However, the central idea is that some portion of a physical universe, including some conscious reasoning agents that exist in that cosmological universe, might in fact be part of a simulation being run in a physical computer of some different, super-sophisticated race of aliens [12, 31, 20]. Under the simulation hypothesis, those “conscious reasoning agents” would just be variables evolving in a computer program running in that physical computer designed by that putative super-sophisticated race of aliens.
In particular, it might be that a portion of our physical universe is being simulated this way, and that in fact we are just variables evolving in some computer in that physical universe. The vague implication is that if we are in fact a simulation, then what we perceive as reality does not have any sort of “objective truth”, since it is just a simulation. What we think of as “reality” would in fact be a chimera, simply a play being put on by the super-sophisticated race of aliens. Crucially, that play could just as easily be put on by some other super-sophisticated race of aliens, in some other physical universe — what we perceive as reality is just the play, not the physical universe in which the play is being staged.
Much of philosophy of science, going back to Kant (at least), would be rendered moot if we are simulations. In particular, this would be the case for many of the flavors of “*** realism” (e.g., scientific realism), which suppose there “exists” some “real” physical truth which is not reducible to a mathematical formalism, a physical truth that instead is somehow “concrete”, and non-mathematical in its very essence — and so not just a computer algorithm. [^1] Arguably, the simple fact that the simulation hypothesis is logically consistent, but if true would imply that there is no such “concrete”, non-mathematical reality, establishes that it is impossible to establish the logical necessity of any of these flavors of realism,
@@
These investigations of relative probabilities all beg (many) questions, about what such probabilities might actually mean, formally. (Just as there are such questions concerning the idea of ascribing probabilities to the universes in a level IV multiverse in general.) How does one ascribe a measure, obeying the Kolmogorov axioms based on some associated sigma algebra, to a collection of events each of which is a universe? Note that it is not even clear that this collection would be a set, rather than a proper class of some sort. Directly reflecting such problems, one could not experimentally assess any proposed value of such a probability, e.g., with a proper scoring rule. [^2] One might try to use a Bayesian “degree of belief” interpretation of probability to circumvent this issue. However, without any decision that would be made based on the probability assigned, and associated loss function, one could not apply Bayesian decision theory. So it is hard to see how “degree of belief” is meaningful in this case. See [62] for some related arguments.
More prosaically, many papers considering such probabilities assume, implicitly or otherwise, that there is some way to assign a uniform probability to all universes. It is proven below in Section 2.4 that this is mathematically those in the kind of universes considered in this paper (which, one might argue, is the kind of universes considered in these other papers). One might imagine that rather than a uniform probability distribution, one could assign some sort of Cantor (“fair coin”) measure to the set of possible universes, as is done for example in algorithmic information theory. However, it is proven in Section 7 that it is impossible to assign a Cantor measure to a set of universes obeying a natural set of restrictions.
@@
Yet others have considered issues concerning the simulation hypothesis that might charitably be characterized as “philosophical” in nature [32, 21, 26, 37], at least in that they do not involve mathematical analysis of any sort.
@@
One earlier set of semi-formal results along these lines was presented in [65]. Some of those semi-formal results define “simulation” in terms of the relationship between languages at different levels of the time, polynomial, or exponential hierarchies of CS theory, and some of them define “simulation” in terms of the relationship between languages at different levels the arithmetic or analytic hierarchies of logic, or similar constructions. (See [22, 5, 30] for background on such hierarchies.) Other results in that paper define “simulation” in terms of the relationship between computational machines with different Turing degrees [58, 55]. Yet others consider the application of Gödel’s second incompleteness theorem, defining simulation in terms of languages (in the logic theory sense) with nested sets of axioms.
None of these earlier results in the literature consider what it means for a computational machine to simulate a physical universe, per se. There is no concern for “coupling” the mathematics of CS theory to the laws of physics of our universe. One natural way to address this lacuna is to use a particular formalization of the physical Church-Turing thesis (PCT), and a closely related thesis which I call the reverse PCT (RPCT). That is the approach I adopt in this paper. [^3] The reader unfamiliar with TMs and associated concepts like Universal Turing machines (UTMs), instantaneous descriptions (ID) of the state of a TM and its tapes, prefix-free encodings, etc., should consult Appendix A for appropriate background. Specifically, in this paper I will say that the PCT applies to a particular (portion of a) universe if the dynamics of that (portion of a) universe can be implemented on a Turing machine (TM). The RPCT instead says that any desired TM can be implemented by the dynamics of the universe in question, by appropriate choice of the initial conditions of that universe.
@@ 2.1 Notation
Other important definitions and notations is in Appendix A , e.g., of partial functions, and of computable functions (here always assumed to be total). I also review the definition of “computational universality” in that appendix. In addition, there I review the definitions of a universal TM (UTM), and a prefix-free TM (the implementation of TMs I will often be assume in this paper, implicitly or otherwise). I also define the instantaneous description (ID) of a Turing machine (TM) there. In the main body of this paper I will assume that the non-blank alphabet of the TM, $\Lambda$, is just ${\mathbb{B}}$. So as described in Appendix A, we can take the state of the tape to be a finite string in ${\mathbb{B}}^{*}$, even though strictly speaking the definition of TMs assumes infinitely long tapes.
@@ 2.2 Framework for analyzing CS theory of the simulation hypothesis
where $W$ is initially finite and cannot be extended dynamically, whereas $N$ is countably infinite. I will parameterize $N$ as $\underline{N}\times R$.
In this decomposition, $R$ is a finite set that represents the internal variables of the computer. So for example, in a conventionally represented TM, $R$ would be the state of the TM’s head, the position(s) of its pointer(s) on the tape(s), etc.  [^5] Strictly speaking, in a TM the pointer variables can have any value in ${\mathbb{N}}$, since they can point to an arbitrary position on the (countably infinite) tape(s). This technical concern can be addressed several ways. For example, we might want to have the states of the pointer variables at any iteration after the $k$’th iteration given by the state of some special subsection of the tape(s), $\underline{n}$, with the states of those variables at earlier iterations recorded in $R$. Another possibility would be to have $R$ infinitely extendable.
$\underline{N}={\mathbb{B}}^{*}$ is instead the state of the tape(s) in the case that the computer is represented as a TM, or it is the state of the memory in the case that the computer is represented as a RAM machine, or an appropriately modified laptop computer. The reason for writing the countably infinite space $\underline{N}$ as the set of finite bit strings is so that we can physically implement it as a system that initially has a finite state space, but whose state space can be expanded by an arbitrarily large finite amount an arbitrary number of times as the universe evolves. See Section 2.6. I will write the elements of $\underline{N}$ as $\underline{n}$.
@@
To simplify the analysis, I do not assume that the models of physical universes that I investigate in this paper capture those universes in toto, e.g., if the state spaces of those entire universes are actually uncountably infinite. Similarly, I will not assume that a simulation running in a computer inside a universe models the dynamics of an entire universe. Rather throughout this paper the expression “universe” should be understand as shorthand for “possibly coarse-grained subset of a universe”. However I do assume that the universe obeys deterministic dynamics, be it over the original state space (assuming it is countable) or some coarse-grained state space. As an example, in Section 2.6 I present a detailed example of what such a “coarse-grained subset of a universe” could be for our particular physical universe.
@@ 2.3 What it means for one universe to simulate another
In this paper $V$ “simulates” $V^{\prime}$ if for all initial states $v^{\prime}_{0}$ of $V$, and all amounts of time $\Delta t^{\prime}$ into the future of $V^{\prime}$ that we might to simulate the dynamics of $V^{\prime}$, there is some associated initial state of $N$ and of $W$, together with a time into the future of the simulating universe, $\Delta t$, such that the state of the computer $N$ at that future time $\Delta t$ gives us the desired state of $V^{\prime}$ at its future time $\Delta t^{\prime}$.
I formalize this as follows:
@@
Note that the LHS of Eq. 3 is the second of the two components of the vector-valued function $g$, while the RHS is an encoding of both components of $g^{\prime}$ into a single variable. Note also the requirement in Definition 1 that Eq. 3 hold for all $t\geq\tau$, and so the state of the simulating computer $N$ does not change after it completes its simulation of the future state of $V^{\prime}$. This just means that I require that the simulating computer halts when it completes its simulation. I also require that the ID of the computer $N$ be replaced in toto (i.e., uniquely, with all other variables fixed to some predefined values) by the output of its simulation program. So for example, if $N$ is a multi-tape prefix TM, this means that when $N$ halts with the result of its simulation on its output tape, all the other tapes — the intermediate work tapes and the input tape — have been re-initialized to be all blanks.
@@
Note that Definition 1 does not require that the simulating computer $N$ calculates the future state of the universe $V^{\prime}$ by itself, independently of the initial state of the environment $W$ outside of $N$, i.e., independently of the value of $w_{0}$. (Formally, the LHS of Definition 1 is not independent of $\omega$, the initial state of $W$.) The reason for this flexibility is to allow the computer $N$ to retrieve the specific information it needs to perform its simulation of the dynamics of the specific state $v^{\prime}_{0}$ from those super-aliens who are running that computer $N$, and who exist in the environment of $N$, $W$. Note though that this freedom also allows the beings running the simulation computer $N$ to intervene on the dynamics of that computer, e.g., by overwriting the simulation program being run in the computer at any time they want. They can even “pull the plug early” on that computer, before it finishes its computation.
@@
I will informally use the term cosmological universe to refer to an entire physical universe obeying one set of laws of physics throughout, with one set of shared initial conditions, etc. [^6] For simplicity, I sidestep the issue of multiverses with different physical constants but otherwise identical physical laws, e.g., arising from a shared inflation epoch. In general, $V$ and $V^{\prime}$ might be portions of different cosmological universes, obeying different laws of physics. They could also be sub-regions of the same cosmological universe though (and therefore obey the same laws of physics). One way this could occur is if there are physically distinct regions of same universe. (See Section 2.6.) This case of multiple computational universes in the same cosmological universe will be important in the discussion of the simulation graph below in Section 7.
@@ 2.4 The Physical Church-Turing Thesis
Much of the earlier literature on the “physical Church-Turing thesis” accords a prominent role to humans, and their abilities (or lack thereof), e.g., in arbitrarily configuring the initial state of physical systems, or in observing their subsequent physical state. There is no such role in Definition 3. All the PCT means in this paper is that evolving the computational universe $V$ does not require a computational machine more powerful than a TM. In fact, it could be that that evolution can be calculated on a machine that is strictly weaker than a TM, e.g., a finite automaton. Furthermore, if $V$ is a sub-region of some cosmological universe, it could be that it requires machines more powerful than a TM to calculate the evolution of some other sub-region of that cosmological universe, different from $V$.
For these kinds of reasons, some readers might argue that Definition 3 doesn’t exactly capture any of the various properties that have been referred to as the “physical Church Turing thesis” in the literature. In some senses it more like one of the related concepts inspired by modern physics, e.g., some forms of ontic structural realism [28, 40, 39, 4] or the level IV multiverse [59, 60]. But for current purposes, we can ignore these semantic distinctions.
@@ 2.5 The Reverse Physical Church-Turing Thesis
Broadly speaking, the RPCT says that the system $N$ operates like a UTM for all pairs $(k,y)$ such that $T^{k}(y)$ is defined, where $k$ and/or $y$ may be encoded in some degrees of freedom in $w_{0}$ rather than directly in $n_{0}$. (That freedom to have $k$ and / or $y$ specified in the environment of the computer $N$ allows that computer to do things like observe its environment to retrieve the input string for a computation from its environment before running that computation.)
@@
Eqs. 6 and 7 mean that the physical system implementing the UTM is initialized with the precise TM that it is to supposed to implement, but not the actual data that TM will be run on. In general, that data (the value of $y$) would be transferred into the computer $N$ at a subsequent iteration, after this initialization of the UTM, but before the UTM starts to run. As an example, this is what would occur if we were to initialize a laptop with a universe-simulating program, with the precise data that program is to run on fed into the laptop before starting the simulation program. (See Section 2.6.) In this paper, when assuming the RPCT I will not implicitly assume that the pristine RPCT holds unless I explicitly say so.
Just as the PCT as defined in Definition 3 has no role for humans, the RPCT has no role for them. In particular, Definition 4 does not say that humans could configure the initial state of $V$ to implement the dynamics of any desired TM. It simply says that there is some such initial state of $V$ that could implement that dynamics.
@@
Despite the popularity of the RPCT among researchers, even if the PCT holds in our specific universe, that does not mean that the RPCT has to hold as well. In particular, cosmological considerations could prevent it from holding [38]. So the results below specific to universes that obey the PCT and / or universes that obey the RPCT might not apply to our specific universe. However, approximations of the analysis might apply even if the PCT and/or RPCT don’t hold exactly in our universe, depending on how precisely they are violated.